if two charges having masses m and 2m are placed in electric field having charges 4q and 2q what is ratio of kinetic energy

Grade:12

2 Answers

Arun
6 years ago
let E  = uniform electric field
let m1 = m
m2 = 2m
force on particles 
F1 = m a1 = 4q E
F2 = 2m a2 = 2q E
a1 = 4q E/m
a2 = 2q E/2m = q E/m
 final velocities of particle
v1 = a1 t = 4q E* t/m
v2 = a2 t = q E* t/m
hence K1 = (1/2) m * 16q2 E2 t2/ m2
K2 = (1/2) 2m * q2 E2 t2/ m2
so
K1 / K2 = 8/1
Chandan
6 years ago
ke=kcharge
ke=k/mass
where k= constant
therefore ke=1/8
hope it helps please approve if it does